OpenStudy (anonymous):

Solve the equation Quantity x plus 1 all over 3 = 2 for x. Please show your work and solution.

OpenStudy (ivettef365):

so you have x+1 ---- = 2 3

OpenStudy (anonymous):

\[\frac{x+1}{3}=2\]??

OpenStudy (hba):

Follow these steps: 1) Multiply both sides by 3 2)subtract both sides by 1

OpenStudy (anonymous):

1) multiply both sides by 3 ) subtract 1 from both sides

OpenStudy (anonymous):

idk how to do that

OpenStudy (anonymous):

i don't believe you what is \(3\times 2\) ?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

\[\frac{x+1}{3}=2\] \[\frac{x+1}{3}\times 3=2\times 3\]\[\frac{x+1}{\cancel{3}}\times \cancel{3}=2\times 3\] \[x+1=6\]

OpenStudy (anonymous):

i figured you could multiply 2 times 3 and get 6, maybe what wasn't clear is if you multiply the left side by 3, the denominator goes, leaving only \(x+1\) on the left

OpenStudy (anonymous):

so that's the answer?

OpenStudy (anonymous):

not quite, almost though what number would you add one to in order to get 6 ?

OpenStudy (anonymous):

Hope this will let you understand what you should be doing on this problem .. ( X + 1) / 3 = 2 2. Multiply both sides of the equation by 3 [(X+1)/3] * 3 = 2 * 3 3/3 = 1, and 1 times anything is that thing itself, so (X+1) * 3 /3 = X+1 And 2 * 3 = 6, of course. X + 1 = 6 3. Subtract 1 from both sides to give the variable, X, by itself X+1 - 1 = 6 - 1 X + 0 = 5 X = 5

OpenStudy (anonymous):

@nunncm ??do you understand know ?

OpenStudy (anonymous):

yeah thankss

OpenStudy (anonymous):

could you give me a best response star pleas @nunncm

OpenStudy (anonymous):

thanks